# Extracting the singularity of a logarithmic partition function to investigate quantum phase transitions

### v2

@@ Abstract
For the O(3) QCP, fitting the RTE data yields $J_{1,c}=0.523(1)$, $\nu=0.70(1)$, while the DRTE gives $J_{1,c}=0.525(1)$, $\nu=0.71(2)$. These values are consistent with the known critical parameters $J_{1,c}=0.52337$ and $\nu=0.7073$ [44, 46, 47]. The universal scaling functions near the O(2) and O(3) QCPs, obtained from data collapse, are shown in Figs. 1(f,h) and (j,m), respectively.
First-order and weakly first-order transitions.— Having established the universal scaling behavior at (2+1)D QCPs, we now turn to first-order phase transitions, with a focus on weakly first-order transitions, which presents significant challenges for traditional local physical quantities.
We first study the 2D checkerboard $J$-$Q$ model ( see SM for model details), which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$ from the O(3)-symmetry breaking AF order to the $\mathbb{Z}_{2}$-symmetry breaking plaquette-singlet solid order [49]. The DRTE around this transition is plotted in Fig. 2 (a). A clear discontinuity in DRTE is observed at the phase transition point, with the DRTE diverging toward opposite signs on either side as $L\to\infty$. This behavior arises because DRTE isolates the singular contribution to the logarithmic partition function (as mentioned above). For comparison, we also calculate the derivative of logarithmic partition functions, as shown in Figs. 2 (b) and (c). These curves appear much smoother across the transition compared with DRTE, due to the dominance of a non-singular, volume-law term. The phase transition of this model is obviously first-order according to the previous works [49, 29], thus we see that the curves in (b) and (c) have very small kinks around the phase transition point, especially for larger sizes.
@@
The phase transition between AF Néel order and valence-bond solid (VBS) phase in 2D quantum magnets can appear (proximately) continuous, which is thought as the deconfined quantum criticality (DQC) [4, 5]. $J-Q_{2}$ and $J-Q_{3}$ models are prototypical examples that realize such transitions in 2D quantum spin systems [50, 51]. These models are free of the sign problem, making them ideal for QMC simulations. Extensive QMC studies have shown that the phase transitions of the JQJ-Q models on the 2D lattice are continuous and exhibit emergent SO(5) symmetry, which is a hallmark of DQC [50, 51, 52, 53, 54, 55]. However, the critical exponents observed in these models do not satisfy the conformal bootstrap bounds expected for SO(5) critical point [56]. Recent works have increasingly suggested that this DQC scenario may instead correspond to a weakly first-order transition or as a nonunitary CFT  [7, 11, 57, 58, 12, 59, 60, 61, 8, 15, 62, 63, 64, 65, 16]. In the following, we show that the DRTE provides a clear and direct signature that the transitions in the $J-Q_{2}$ and $J-Q_{3}$ models ( see SM for models’ details) are indeed first-order, even in rather small systems (up to $L=56$).
@@
The transition points of the 2D square lattice $J-Q_{2}$ and $J-Q_{3}$ models have been numerically determined as $J_{c}/Q=0.04502$ and $J_{c}/Q=0.67046$, respectively [50, 51, 66, 63]. The DRTE results for both models are shown in Figs. 2 (d) and (g), along with the derivatives of the logarithmic partition function in panels (e,f) and (h,i). Due to the dominant background contribution, the derivative of the logarithmic partition function appears smooth across the transition, even for large system sizes, as seen in Figs. 2(e,f) for $J-Q_{3}$ and (h,i) for $J-Q_{2}$ model. This shows that the derivative of the logarithmic partition function can not be directly reveal the weakly first-order nature. In contrast, the DRTE reveals clear discontinuities at the transition points, providing direct evidence that both transitions are (weakly) first-order. Notably, the $J-Q_{2}$ model exhibits stronger finite-size effects than $J-Q_{3}$, which has made its first-order nature more difficult to resolve in previous studies [63]. Our results demonstrate that DRTE can unambiguously detect even weakly first-order transitions, in relatively small systems (up to $L=56$).
Conclusion and discussions.— We proposed that the RTE isolates the singular contribution of the logarithmic partition function, making it highly sensitive near phase transitions. We systematically studied the RTE and its derivative DRTE around QCPs in 2D quantum spin systems. By deriving universal scaling forms, we successfully captured the numerical data and extracted the critical exponent $\nu$ via data collapse near (2+1)D QCPs. Importantly, by definition, the DRTE provides a clear signature to distinguish weakly first-order phase transitions and deconfined quantum critical points even in small sizes, which is highly challenging to detect using conventional local observables.
Notably, RTE and DRTE are universally applicable, as their definitions do not depend on specific models or order parameters. They also have the potential to characterizing phase transitions that are difficult to probe with conventional local observables, such as topological transitions, and may reveal new universal critical behaviors. Our work establishes a general and effective framework for studying quantum phase transitions across a wide range of classes, including continuous, first-order, and particularly elusive weakly first-order transitions.
@@ Acknowledgements.
We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang and Wenan Guo. Z.W. thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. The work is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
Supplemental Materials
@@ Appendix A Method
In this frame, we can designed the incremental process along the path composed of real physical parameters  [^1] It was virtual path before. In other words, all intermediates are physical partition functions at corresponding parameters. The continuously incremental trick here are introduce as follows:

### v3

@@ II.4 Continuous symmetry phase transitions
For the O(3) QCP, fitting the RTE data yields $J_{1,c}=0.523(1)$, $\nu=0.703(7)$, while the DRTE gives $J_{1,c}=0.525(1)$, $\nu=0.708(8)$. These values are consistent with the known critical parameters $J_{1,c}=0.52337$ and $\nu=0.7073$ [63, 65, 66]. The universal scaling functions near the O(2) and O(3) QCPs, obtained from data collapse, are shown in Figs. 1(f,h) and (j,m), respectively.
In summary, all the data show that the RTE does have a crossing at a critical point and the DRTE does extract the critical exponent $\nu$ by data collapse, which is highly consistent with our theoretical analysis.
III First-order phase transition
III.1 Finite-size scaling form
At a first-order phase transition point $g_{c}$, the free energy densities of the two phases, $f_{1}(g)$ and $f_{2}(g)$, are equal, but their slopes differ (i.e., $\Delta f^{\prime}(g_{c})\neq 0$). Near the first-order phase transition point, for a finite system of linear size $L$ and spatial dimension $d$, at sufficiently low temperatures, the partition function of the system can be approximated as a simple sum of contributions from the two phases [70]: $Z(g,\beta,L)\simeq e^{-\beta L^{d}f_{1}(g)}+e^{-\beta L^{d}f_{2}(g)}.$ Here, contributions from the interface between the two phases and corrections due to fluctuations are neglected; this approximation holds when $L$ is large and the temperature is low. Each term corresponds to the free energy of a pure phase, $F_{i}=L^{d}f_{i}(g)$.
The logarithm of the partition function can be expressed as $\ln Z(g,\beta,L)=-\beta L^{d}f_{1}(g)+\ln\!\left[1+e^{-\beta L^{d}\Delta f(g)}\right]$, where we define the free energy density difference as $\Delta f(g)=f_{2}(g)-f_{1}(g)$. From this expression, we can further derive the second Rényi entropy as
$S^{(2)}=2\ln\!\left(1+e^{-\beta L^{d}\Delta f}\right)-\ln\!\left(1+e^{-2\beta L^{d}\Delta f}\right).$ (11)
Thus, the DRTE takes the compact expression
$\frac{\partial S^{(2)}}{\partial g}=2\beta L^{d}\,\Delta f^{\prime}(g)\bigl[w(2\beta)-w(\beta)\bigr],$ (12)
where $w(\beta)=\frac{1}{1+e^{-\beta L^{d}\Delta f(g)}}$ is the logistic function, which here represents the statistical weight associated with one of the two phases.
Figure 2: (a) Derivative of the RTE, and (b) corresponding data collapse analysis near the phase transition point of the 2D anisotropic Heisenberg model with anisotropy parameter $\Delta$ ( see Eq. (10)). (c) and (d) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition. Here, $J_{1}=J_{2}=1$, and the anisotropy $\Delta$ is taken as the tuning parameter. (e–h) Results for the checkerboard $J$-$Q$ model, using $Q=1$ and tuning $J$.
The Eq. (12) immediately explains 3 characteristic features of the DRTE at a first-order transition:
(1) Since $\Delta f^{\prime}(g_{c})\neq 0$, the difference $w(2\beta)-w(\beta)$ changes sign across $g_{c}$ (near $g_{c}$, $\Delta f(g)\approx\Delta f^{\prime}(g_{c})(g-g_{c})$, crossing zero linearly), resulting in a robust double-peak structure with one positive and one negative peak located on opposite sides of the transition. This discontinuous structure reflects the rapid transfer of thermodynamic weight between the two competing phases and is a direct consequence of phase coexistence. Indeed, this is fully consistent with the definition of a first-order phase transition. The RTE is strictly defined as the difference in the logarithm of the partition function at different inverse temperatures: $S^{(n)}=n\ln Z(\beta)-\ln Z(n\beta)$, where different values of $n$ correspond to different linear rates of approaching zero temperature. For quantum phase transitions, the inverse temperature $\beta$ is a non-relevant field, and thus the difference in the logarithm of the partition function corresponds to the difference in free energy. Mathematically, the sum or difference of two continuous functions remains continuous. In other words, if the resulting function is discontinuous, then at least one of the original functions must be discontinuous. In this case, the DRTE is such a resultant function, and its discontinuity directly reflects a discontinuity in the first derivative of the free energy itself. Therefore, the discontinuity in the DRTE directly characterizes a first-order phase transition.
(2) Due to $\Delta f(g_{c})=f_{2}(g_{c})-f_{1}(g_{c})=0$, it follows that $w(2\beta)-w(\beta)=0$, and therefore the DRTE equals zero at the transition point. This represents another fundamental distinction from continuous phase transitions. In continuous transitions, the divergence of the correlation length leads to a critical singularity, resulting in a divergent behavior of the DRTE that scales with system size as $L^{1/\nu}$ (as shown in Eq. (7)).
(3) Because $|w(2\beta)-w(\beta)|$ is an $O(1)$ constant, and $\Delta f^{\prime}(g)$ is also of constant order, it follows that the peak amplitude of the DRTE scales as
$\left|\partial S^{(2)}/\partial g\right|\sim\beta L^{d}\sim L^{d+z}.$ (13)
For quantum phase transitions in $(2+1)$ dimensions with dynamical exponent $z=1$ (so that $\beta\sim L$), gives $\left|\partial S^{(2)}/\partial g\right|\sim L^{3}.$ For a first-order phase transition, the inverse of the correlation length exponent is given by $1/\nu=d+z$ effectively (here, $1/\nu=3$). It implies that DRTE satisfies the following finite-size scaling ansatz:
$\left|\partial S^{(2)}/\partial g\right|=L^{d+z}f(\Delta gL^{d+z}).$ (14)
It is worth noting that the DRTE characterizes the difference in free energy, rather than the difference in free energy density. This divergence does not arise from a critical singularity, but rather stems from the intrinsic volume-law scaling of the system and the fact that the free energies of the coexisting phases are no longer equal once the system moves away from the transition point.
III.2 Numerical demonstration
Our numerical simulations confirm our theoretical derivation. Specifically, we study the 2D anisotropic Heisenberg model and the 2D checkerboard $J$-$Q$ (CBJQ) model [71]. Let us begin by examining the 2D anisotropic Heisenberg model defined in Eq. (10)), where the antiferromagnetic couplings are uniform, $J_{1}=J_{2}=1$, and the anisotropy parameter $\Delta$ serves as the tuning variable. At $\Delta=1$, the system exhibits full spin-rotation symmetry, corresponding to O(3), which is spontaneously broken by the emergence of AF order in the ground state. When $\Delta$ deviates from unity, the symmetry is explicitly lowered to in-plane spin rotation symmetry and spin-flip symmetry. For $0<\Delta<1$, the ground state develops an easy-plane AF order, spontaneously breaking the continuous $\mathrm{SO}(2)$ symmetry. In contrast, for $\Delta>1$, the system favors an easy-axis AF order, leading to the breaking of the discrete $\mathbb{Z}_{2}$ spin-inversion symmetry. Precisely at $\Delta=1$, both types of order coexist, and each vanishes discontinuously across the transition. This abrupt change in the order parameters signals a first-order quantum phase transition, accompanied by the spontaneous breaking of an enhanced O(3) symmetry [71].
The DRTE across the transition at $\Delta_{c}=1$ is shown in Fig. 2(a). A clear discontinuity is observed near the phase transition point, with the DRTE diverging toward opposite signs on either side of the transition in the thermodynamic limit ($L\to\infty$). For comparison, we also compute the derivative of the logarithm of the partition function, as displayed in Figs. 2(c) and (d). These curves exhibit a significantly smoother behavior across the transition compared to the DRTE. Moreover, the DRTE curves for different system sizes are seen to intersect precisely at the critical point $\Delta_{c}=1$, and the value at the crossing point is zero—consistent with the expected vanishing of the DRTE at a first-order transition. Finally, as shown in the data collapse in Fig. 2(b), the DRTE obeys the scaling form given by Eq. (14). These results provide strong numerical confirmation of our theoretical predictions.
The occurrence of a symmetry-enhanced first-order transition is not limited to systems with exact symmetries. An emergent enhanced symmetry can arise at long wavelengths provided that all explicit symmetry-breaking perturbations are irrelevant under the renormalization group (RG) and thus decay along the RG flow. In this scenario, we study the 2D checkerboard $J$-$Q$ model (see Appendix B for model details), which exhibits a spontaneously broken emergent O(4) symmetry at the first-order transition point $J_{c}/Q=0.217(1)$, where the system transitions from an O(3)-symmetry-breaking antiferromagnetic order to a $\mathbb{Z}_{2}$-symmetry-breaking plaquette-singlet solid order [71]. The DRTE across this transition is shown in Fig. 2(e). A clear discontinuity is observed at the critical point, with the DRTE diverging to opposite signs on either side of the transition in the thermodynamic limit ($L\to\infty$). The RTE equals zero at the transition point. Furthermore, the DRTE data collapse successfully onto a single curve when rescaled according to Eq. (14) (Fig. 2(f)), consistently confirming the validity of our theoretical predictions. Similar to the anisotropic Heisenberg model, the derivative of the logarithm of the partition function (shown in Figs. 2(g) and (h)) exhibits a much smoother behavior across the transition compared to the DRTE.
Combining theoretical analysis with numerical simulations, we demonstrate that the DRTE is not only effective for characterizing continuous phase transitions, but also serves as a powerful probe for first-order phase transitions.
IV weak first-order transitions
IV.1 Finite-size scaling form
Having established the universal behaviors at second order QCPs and first order phase transitions, we now turn our attention to weak first order transitions, which pose significant challenges for conventional local physical observables. The fundamental challenge is that there is a very large correlation length at the phase transition point, thus all the behaviors in finite sizes are similar to a critical point.
The weak first order phase transitions are generally believed to be governed by two nearby fixed points in finite sizes: one corresponding to the first order transition that the system flows into in the thermodynamic limit, and the other associated with a continuous phase transition in close proximity. For a continuous transition, the correlation length diverges near the critical point as $\xi\sim|J-J_{c}|^{-\nu}$. When these two fixed points are very close—i.e., when $|J-J_{c}|$ is small—the correlation length $\xi$ can become large even for a slight deviation from criticality. As a result, for finite system sizes with $L<<\xi$, the system lies within the critical region of the nearby continuous transition, even as it approaches the first-order transition point. Consequently, the system exhibits signatures of continuous critical behavior, despite the underlying transition being first-order. This crossover effect explains the observed scaling collapse using continuous-transition exponents in finite-size systems.
Figure 3: We set $Q=1$ and use $J$ as the tuning parameter. (a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$. (b) Corresponding data collapse analysis of (a). (c) and (d) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the $J-Q_{3}$ model. (e–h) Data for the $J-Q_{2}$ model. The raw values of the partition function derivative vary significantly across system sizes (see Appendix B), making direct comparison difficult. To resolve this, we normalize the data by $L^{3}$, corresponding to $\beta L^{2}=L^{3}$.
Let us go back to the scaling behavior of DRTE that captures characteristics of the standard first order transition in Eq. (12), where actually we have ignored the quantum fluctuations since it is a strong first order phase transition. However, in the case of a weak first-order transition, quantum fluctuations become essential and must be properly accounted for. In finite-size simulation, quantum fluctuations originating from a nearby critical point introduce an effective “singular” contribution—though not a true thermodynamic singularity—that mimics the behavior observed in second-order transitions (as shown in Eq.(7)). We therefore conjecture that the DRTE expression in Eq.(12) should be modified in the vicinity of such a critical point to include this effect:
$\frac{\partial S^{(2)}}{\partial g}=(2L^{d+z}\,\Delta f^{\prime}(g)+\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu})\bigl[w(2\beta)-w(\beta)\bigr],$ (15)
The Eq. (15) predicts some features of DRTE behaviors at a weak first order phase transition point:
1) The double-peak structure in first order phase transition is still present, since the $w(2\beta)-w(\beta)$ term remains.
2) The DRTE also equals zero at the transition point, since the fact of energy-crossing at first order phase transition is unchanged.
3) The DRTE does not follow the finite-size scaling form for a strong first-order transition in Eq. (14) for system sizes smaller than the correlation length, but instead obeys the “singular” correction term scaling as $\sim L^{1/\nu}$, where $\nu$ reflects the critical exponent of the nearby critical point. For the part $(2L^{d+z}\,\Delta f^{\prime}(g)+\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu})$ in Eq.(15), we argue that the coefficient of the term $2L^{d+z}\Delta f^{\prime}(g)$, i.e., $2\Delta f^{\prime}(g)$ is extremely small around the weak first order phase transition point, thus it will suppress the $L^{d+z}$ scaling in finite sizes and emphasize the $\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}$ scaling term. While the simulating sizes become much larger than the correlation length, the $L^{d+z}$ scaling dominates and the “singular" term can be ignored due to $d+z>1/\nu$, the Eq.(15) goes back to Eq.(12).
In the following, we will show the numerical data which strongly support our analysis above.
IV.2 Numerical demonstration
@@
The phase transition between AF Néel order and VBS phase in 2D quantum magnets can appear (approximately) continuous, which is thought as the deconfined quantum criticality (DQC) [72, 73]. $J-Q_{2}$ and $J-Q_{3}$ models are prototypical examples that realize such transitions in 2D quantum spin systems [5, 22]. These models are free of the sign problem, making them ideal for QMC simulations. Extensive QMC studies have shown that the phase transitions of the JQJ-Q models on the 2D lattice are continuous and exhibit emergent $SO(5)$ symmetry, which is a hallmark of DQC [5, 22, 74, 75, 9, 76]. However, the critical exponents observed in these models do not satisfy the conformal bootstrap bounds expected for SO(5) critical point [77]. Recent works have increasingly suggested that this DQC scenario may instead correspond to a weak first order transition  [78, 79, 80, 81, 82, 21, 18, 31, 83, 84, 85]. Next, we will show that the DRTE behaviors provide a clear and direct signature that the transitions in the $J-Q_{2}$ and $J-Q_{3}$ models ( see Appendix B for models’ details) are indeed weak first order in finite sizes (up to $L=56$). The observed DRTE signals fully align with the three key features predicted in the previous subsection, offering compelling evidence against a genuine continuous transition.
@@
The transition points of the 2D square lattice $J-Q_{2}$ and $J-Q_{3}$ models have been numerically determined as $J_{c}/Q=0.04502$ and $J_{c}/Q=0.67046$, respectively [5, 22, 86, 31]. The DRTE results for both models are shown in Figs. 3 (a) and (e), along with the derivatives of the logarithmic partition function in panels (c,d) and (g,h). The derivative of the logarithmic partition function appears smooth across the transition, even for large system sizes, as seen in Figs. 3(c,d) for $J-Q_{3}$ and (g,h) for $J-Q_{2}$ model. This shows that the almost smooth derivative of the logarithmic partition function cannot directly reveal the weak first-order nature, as discussed above, since the logarithmic partition function contains large analytic contributions.
In contrast, the DRTE reveals clear discontinuities with a double-peak character near the transition points (see Figs. 3 (a) and (e)), as the feature 1) we predicted in the above subsection. Moreover, DRTE curves for different system sizes intersect at zero at the transition point (as the feature 2) we predicted above), indicating the absence of singular behavior—unlike in continuous phase transitions. These two features are fully consistent with our previous analysis of the DRTE at weak and standard first order transitions. Our results demonstrate that the DRTE can unambiguously detect the first-order nature of weak first-order phase transitions in relatively small systems (up to $L=56$), even in the $J$-$Q_{2}$ model with much stronger finite-size effects. In fact, its nature, whether continuous or first-order, remains under debate so far [31].
[…]
It is worth noting that, in current studies based on local order parameters, the observed behavior within accessible system sizes is predominantly governed by signatures of continuous phase transitions, making it difficult to unambiguously identify the underlying first-order nature. In contrast, our approach—using the DRTE—successfully captures features of both continuous critical fluctuations and first-order discontinuity even in relatively small systems (up to $L=56$). This provides a clear and compelling demonstration of the weak first-order character of the transition, marking, to our knowledge, the first such identification achieved at modest system sizes.
V Conclusion and discussions
We have established a unified and comprehensive scaling theory for the Rényi thermal entropy (RTE) and its derivative (DRTE) across all three fundamental types of phase transitions: continuous (second-order), strong first-order, and weak first-order. Our framework is predicated on the key insight that the DRTE naturally isolates the singular part of the free energy by canceling dominant analytic contributions. Crucially, we have derived distinct and diagnostic scaling behaviors for the DRTE in each regime:
1. Continuous Phase Transitions: The DRTE scales as
$\frac{\partial S^{(2)}}{\partial g}=L^{1/\nu}\,\tilde{S}^{\prime}(\Delta gL^{1/\nu}),$ (16)
where $\nu$ is the correlation length exponent. This allows for precise determination of the critical point $g_{c}$ and exponent $\nu$ via data collapse, as demonstrated for (2+1)D O($N$) universality classes.
2. Standard First-Order Transitions: The DRTE exhibits a characteristic discontinuity with a double-peak structure and follows the volume-law scaling
$\frac{\partial S^{(2)}}{\partial g}=\left[2L^{d+z}\Delta f^{\prime}(g)\right]\left[w(2\beta)-w(\beta)\right]$ (17)
where $d$ is the spatial dimension and $z$ is the dynamical exponent. The zero crossing at $g_{c}$ and the sign change of the peaks provide an unambiguous signature of phase coexistence.
3. Weak First-Order Transitions: The DRTE obeys a hybrid scaling form
$\frac{\partial S^{(2)}}{\partial g}=\left[2L^{d+z}\Delta f^{\prime}(g)+\tilde{S}^{\prime}(\Delta gL^{1/\nu})L^{1/\nu}\right]\left[w(2\beta)-w(\beta)\right],$ (18)
which encapsulates the competition between the underlying first-order mechanism (first term) and the influence of a nearby critical fixed point (second term). In accessible finite sizes ($L\ll\xi$), the critical-like $L^{1/\nu}$ scaling dominates, explaining the observed data collapse with continuous exponents. Simultaneously, the persistent double-peak structure and zero crossing betray the first-order nature, as conclusively shown for the debated $J$-$Q$ models.
Beyond these specific findings, the RTE/DRTE approach offers a general, unbiased, and order-parameter-agnostic diagnostic tool. Its applicability extends to transitions elusive to conventional local observables, including potential topological phase transitions. While validated here with high-precision quantum Monte Carlo, our method is readily adaptable to tensor network and other numerical techniques [87, 88]. This work thus provides a new paradigm for characterizing quantum criticality, resolving the long-standing challenge of distinguishing weak first-order from continuous transitions in finite-size simulations.
@@ Acknowledgements.
We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang, Zi Yang Meng, Chong Wang and Wenan Guo. Z.W. thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. The work is supported by the Scientific Research Project (No.WU2025B011) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
@@ Appendix A Method
In this frame, we can designed the incremental process along the path composed of real physical parameters. In other words, all intermediates are physical partition functions at corresponding parameters. The continuously incremental trick here are introduce as follows: